OpenStudy (anonymous):

Any help would be greatly appreciated :) Partial derivatives: Find the partial derivative of f(x,y) = ln(x+y)/(x-y)

OpenStudy (anonymous):

do each separately.

OpenStudy (anonymous):

Partial derivative of x: take the usual derivative, but treat y as a constant. Partial derivative of y: take the usual derivative, but treat x as a constant. Expression you have is not crystal clear. Is (x-y) within the logarithm argument or dividing ln(x+y)?

OpenStudy (anonymous):

The textbook states it like that but I believe it is ln((x+y)/(x-y))

OpenStudy (anonymous):

ln[(x+y)/(x-y)] = ln(x+y) - ln(x-y) simplifying your life! (d/dx) ln(x+y) = 1/(x+y) etc. (d/dy)lin(x+y) = 1/(x+y) etc.